# A Typology of Decision-Making Tasks for Visualization

### v1

@@ Abstract
Despite decision-making being a vital goal of data visualization, little work has been done to differentiate the decision-making tasks within our field. While visualization task taxonomies and typologies exist, they are often too granular for describing complex decision goals and decision-making processes, thus limiting their potential use in designing decision-support tools. In this paper, we contribute a typology of decision-making tasks that were iteratively refined from a list of design goals distilled from a literature review. Our typology is concise and consists of only three tasks: choose, activate, and create. Originally proposed by the scientific community, we extend and provide definitions for these tasks that are suitable for the visualization community. Our proposed typology offers two benefits. First, it facilitates the composition of decisions using these three tasks, allowing for flexible and clear descriptions across varying complexities and domains. Second, diagrams created using this typology encourage productive discourse between visualization designers and domain experts by abstracting the intricacies of data, thereby promoting clarity and rigorous analysis of decision-making processes. We motivate the use of our typology through four case studies and demonstrate the benefits of our approach through semi-structured interviews conducted with experienced members of the visualization community, comprising academic and industry experts, who have contributed to developing or publishing decision support systems for domain experts. Our interviewees composed diagrams using our typology to delineate the decision-making processes that drive their decision-support tools, demonstrating its descriptive capacity and effectiveness.
@@
decision-making, typology
@@
This figure shows the three decision tasks in our typology: CHOOSE, ACTIVATE, and CREATE. The CHOOSE task assesses options and outputs a subset deemed optimal or best. The ACTIVATE task represents a decision where options are evaluated, and only those that meet or exceed a threshold are returned. The CREATE task represents decisions on assembling, synthesizing, or generating new information. Together, these three decision tasks can be composed to represent complex decision-making problems.
@@ 1 Introduction
Visualization researchers and practitioners alike have suggested that one of the primary goals of visualization systems is to help users make decisions[55, 61, 67, 23]. Many works have come out of the visualization community that have succeeded in helping users make important decisions (e.g., [1, 70, 16, 2, 26, 29], to name a few). However, when discussing these works at the abstract level, the taxonomies within the visualization community (e.g., [3, 74, 13, 23, 24, 48]) do not adequately provide for the complexities embedded within decision-making. Thus, they limit potential use for (1) analyzing and reflecting upon design, (2) helping to facilitate communication and design, and (3) categorizing and organizing publications on decision-making systems.
In this paper, we build upon this past work, most notably Dimara et al.[23, 24, 48, 47], and present a typology of decision-making for visualization. It consists of three decision tasks that represent specific decision problems: CHOOSE, ACTIVATE, and CREATE. The CHOOSE task assesses alternatives and outputs a subset deemed optimal or best, the ACTIVATE task represents a decision where only the inputs that meet or exceed a threshold are returned, and the CREATE task represents decisions that assemble, synthesize, or generate new information. The most critical property of this typology is that it allows us to represent the complex structures of real-world decision-making problems, as our tasks can be composed or decomposed into other tasks.
To inform our typology, we gathered design goals from a thorough literature review of decision-support tools within the visualization community. We built upon previously curated corpora[48, 57], focusing our search on papers with domain expert evaluations. We then followed an inductive coding process to identify each paper’s primary decision-making goals and processes. We then took these codes and distilled the design goals for our decision-making typology, including iterativeness, composability, generalizability, simplicity, and others. From the decision-making theory[28, 58] and our design goals, we worked towards an initial set of tasks. We then used our survey to hone and iterate on our tasks until they could accurately characterize the works included. To illustrate the abovementioned properties, we present four case studies using our typology to describe users’ decisions using decision-support tools.
@@
Overall, the result of our validation can be divided into three broad categories: the correctness of the typology, the utility of the typology, and how the typology compared to existing task taxonomies and typologies. The correctness of the typology can be further broken down into the consideration of completeness and expressiveness. In our interview study, participants found that the typology is complete regarding the three decision tasks and that they could successfully express their decision-making problems in diagrams using the typology. Additionally, participants found the typology to be helpful when analyzing and reflecting on the design process of visualization tools, believed that it could be useful when communicating with folks outside of visualization and that it would be able to categorize visualization decision-making publications. Finally, participants compared the typology to existing work and were excited by our typology’s level of abstraction and flexibility.
@@ 2 Background & Related Work
Decision making is a broad topic studied across many disciplines with visualization researchers emphasizing that supporting decision making is a vital goal of data visualization [23, 35, 4, 15, 5]. The visualization community tends to focus on decision making from three broad lenses: cognition and perception, decision process support, and domain-specific decision support systems. We review the relevant literature from these areas and discuss the significance to this work.
@@ 2.1 Decision Making & Visualization
In a cognitive context, decision-making is often framed in terms of rationality and judgement[51]. In this view, the most rational decisions are ones that maximize expectation of a function[27]. However, a number of human decisions are not rational[60], a fact that has impacted the visualization community in two significant ways. The first is an increase in computational decision-making support[64, 35, 73]. While promising, it is unclear and difficult to determine under what conditions or kinds of decisions users will accept help from or implement decisions of computational models or agents[49, 29] The second impact is the study of cognitive bias and perception in visualization contexts. Cognitive biases explain irrational decision-making as the result of using heuristic functions to evaluate information in a decision-making cognitive process[22].
@@ 2.2 Visualization Typologies
Figure 1: Our processes for: (A) generating a list of design goals from a corpus and (B) using the design goals to create our typology.
@@
Two of the most prominent and influential visualization task typologies in the community are the analytical task typology and the multi-level task typologies. The analytical tasks typology addressed a gap in highly data-centric typologies by taking an analytical-centric approach[3]. This typology consists of ten tasks that describe “primitive" analyses of data, or low-level tasks. The multi-level task typology further expanded upon these tasks to incorporate higher levels that can be composed from lower-level tasks, with the highest-level characterization of tasks falling into three categories of “why?", “how?", and “what?"[44, 13]. However, these tasks are not suitable for describing decision-making in compact yet meaningful ways. Consider the decision of buying a car by choosing among a set of available used cars. This task, in its current form, cannot yet be represented in either the analytical or multi-level task typologies because more information is needed on how the decision is made, which can vary across users. In other words, these typologies are oriented towards describing potential solutions or approaches for solving a problem, they do not describe the decision problem or goal. When we can describe decision problems in abstract yet precise ways, we can begin to reason about solutions more holistically.
@@ 3 Methodology
This section describes our methodology for developing a typology of decision-making tasks. Our approach began with gathering design requirements from a literature review of publications in the visualization community relating to decision-support tools. Specifically, we analyzed design studies and visual analytics application papers to extract example scenarios and relevant contexts relating to decision-making, as detailed in subsection 3.1.
In subsection 3.2, we explain how we formulated these design goals by examining the specific decisions that domain experts sought to achieve within their problem domains while using decision-support tools. This involved dissecting the decision processes, identifying key components, and highlighting other properties relevant to decision-making, as outlined in subsection 3.3.
Following this, we conducted collaborative working sessions with all authors to generate a comprehensive set of decision-making tasks that met the design goals. This was an iterative process that involved continuous refinement and validation of the decision tasks and their definitions, drawing upon prominent decision-making theory papers, as we discuss in subsection 3.4.
The result of this iterative design process is our proposed Typology of Decision-Making Tasks for Visualization, which we present in detail in Section 4.
3.1 Survey Corpus
@@
To refine our selection, we prioritized papers featuring specific domain applications, and eliminated generic systems lacking evaluations with domain experts. This guaranteed that the chosen tools were utilized by domain users with actual decisions and requirements that could be validated. For instance, papers solely based on hypothetical case studies were excluded (e.g., Podium[63], Zooids[42], etc.). This filtering process resulted in a final corpus of 69 papers that directly contribute to our study’s objectives.
@@ 3.2 Coding Process
Our coding process followed an inductive approach to derive the codes, i.e. the design goals for our typology, from the survey corpus outlined in the preceding section. The process involved a team-based approach with two coders, who are the first two authors of this paper. For each paper, we identified the primary decision-making goal by examining the abstract and introduction. We identified a few potential goals and then determined the true goal by analyzing the task abstraction, requirements, evaluation sections, and other supplementary materials, such as figures, video demonstrations, and supplementary materials (if available). In addition, we generated codes for unique decision contexts, sub-decisions that supported the overall decision, and the information that those decisions would use and produce. The resulting codes are discussed in the subsequent subsection.
@@ 3.3 Design Goals for a Decision Task Typology
The generated codes were used to derive a set of design goals. Through multiple brainstorming sessions involving all authors, we distilled the following list of goals that we believe our typology should fulfill in order to accurately represent the decision processes described in the corpus’s case studies and decision-making scenarios. For each goal, we include excerpts from interviews that support our choice of that goal.
@@
    • By working with the domain experts, we constructed a threestage iterative workflow for analyzing and improving bus networks[71]
@@
  • Data-type flexibility: The decision tasks within our typology should demonstrate flexibility and impartiality towards input and output types, instead emphasizing how information is processed and evaluated.
@@
3.4 Iterative Design Process
Our iterative design process began with brainstorming an initial set of decision-making tasks, drawing from the design goals outlined in subsection 3.3. Each iteration involved consulting relevant decision-making theory literature, particularly works by Fischhoff et al.[28, 27], Simon’s decision making phases [58, 14], cognitive science and cognitive process literature[66, 51, 60, 22], to ensure alignment with established theories. In each iteration, we defined the newly introduced tasks and modified or deleted existing task definitions to eliminate overlaps.
Following this, we validated the set of decision tasks by testing them on the survey corpus, aiming to confirm their descriptiveness and properties highlighted in our list of design goals. During this validation phase, we noted any missing capabilities that required inclusion in subsequent iterations. Additionally, we identified tasks that were not utilized for removal in the next iteration.
The iterative process continued with modifications based on feedback from the validation step and revisiting relevant decision-making theory literature to ensure coherence and alignment. This process persisted until the set of decision-making tasks and their definitions reached stability, resulting in a finalized typology, which we describe below.

### v3

@@ Abstract
Despite decision-making being a vital goal of data visualization, little work has been done to differentiate decision-making tasks within the field. While visualization task taxonomies and typologies exist, they often focus on more granular analytical tasks that are too low-level to describe large complex decisions, which can make it difficult to reason about and design decision-support tools. In this paper, we contribute a typology of decision-making tasks that were iteratively refined from a list of design goals distilled from a literature review. Our typology is concise and consists of only three tasks: CHOOSE, ACTIVATE, and CREATE. Although decision types originating in other disciplines exist, we provide definitions for these tasks that are suitable for the visualization community. Our proposed typology offers two benefits. First, the ability to compose and hierarchically organize the tasks enables flexible and clear descriptions of decisions with varying levels of complexities. Second, the typology encourages productive discourse between visualization designers and domain experts by abstracting the intricacies of data, thereby promoting clarity and rigorous analysis of decision-making processes. We demonstrate the benefits of our typology through four case studies, and present an evaluation of the typology from semi-structured interviews with experienced members of the visualization community who have contributed to developing or publishing decision support systems for domain experts. Our interviewees used our typology to delineate the decision-making processes supported by their systems, demonstrating its descriptive capacity and effectiveness. Finally, we present preliminary findings on the usefulness of our typology for visualization design.
@@
decision-making, typology, theory
@@
This figure shows the three decision tasks in our typology: CHOOSE, ACTIVATE, and CREATE. The CHOOSE task assesses options and returns a subset deemed optimal or best. The ACTIVATE task represents a decision where options are evaluated, and only those that meet or exceed a threshold are returned. The CREATE task represents decisions on assembling, synthesizing, or generating new information. Together, these three decision tasks can be composed to represent complex decision-making problems.
@@ 1 Introduction
Visualization researchers and practitioners alike have suggested that one of the primary goals of visualization systems is to help users make decisions[58, 65, 71, 25]. Many works have come out of the visualization community that have succeeded in helping users make important decisions (e.g., [1, 74, 16, 2, 28, 31], to name a few). However, it is difficult to understand patterns across these systems, even among those that address similar decision-making problems in the same domain. Furthermore, the taxonomies within the visualization community (e.g., [3, 78, 13, 25, 26, 51]) do not adequately capture the components of complex decision-making problems and processes because they instead focus on analytical tasks or visual encodings, or they categorize all decision-making tasks under a single “decide” or “choice” task. This leaves a critical gap in understanding how information is processed when making a decision and how individual decisions combine to form a coherent decision-making process.
To address these shortcomings, we introduce a typology specifically focused on decision-making tasks in visualization. It consists of three decision tasks: CHOOSE, ACTIVATE, and CREATE, as shown in Figure A Typology of Decision-Making Tasks for Visualization. Each task assesses options and represents distinct decisions with those options. The CHOOSE task decides which options are optimal or best, the ACTIVATE task decides which options meet or exceed a threshold, and the CREATE task decides how to assemble, synthesize, or generate new information from the options. The most critical properties of this typology are that it allows us to represent the complex structures of real-world decision-making problems through the hierarchical organization of tasks and the composition of tasks.
We designed our typology using design goals derived from a literature review of decision-support tools within the visualization community. We built upon previously curated corpora of application and decision-support visualization systems[51, 60], focusing on those with domain expert evaluations. We then followed an inductive coding process to identify each paper’s primary decision-making goals and processes. From these codes, we distilled the design goals for our decision-making typology, including iterativeness, composability, generalizability, simplicity, and others. From decision-making theory[30, 62] and our design goals, we worked towards an initial set of tasks for the typology. We used our survey to hone and iterate on our tasks until they could accurately characterize the works included. To illustrate that our typology has the aforementioned properties, we present four case studies where authors described the decision-making processes addressed by their tools using our typology.
@@
We divide the results of our validation into three broad categories: the correctness of the typology, the utility of the typology, and how the typology compares to existing task taxonomies and typologies. The correctness of the typology can be further broken down into the consideration of completeness (the typology needs to cover all decision-making problems we have encountered) and expressiveness (the typology can be easily used to describe decision-making scenarios). In our interview study, participants found that the typology is complete regarding the three decision tasks and that they could successfully express their decision-making problems in diagrams using the typology. They also found the typology helpful when analyzing and reflecting on the design process of visualization decision-support tools, believed that it could be useful when communicating with domain experts, and noted that it could categorize visualization decision-making publications. Moreover, participants provided valuable insights when comparing the typology to existing frameworks, highlighting its appropriate level of abstraction for describing decision-making problems and its adaptability to diverse data and contexts.
Lastly, we conducted a preliminary study to evaluate the impact of our typology on visualization design. Our findings revealed that participants used the typology to guide their choices between design alternatives, identify improvements for decision tasks that lacked sufficient support, and even redesign their systems based on the insights gained from applying the typology. With the results from our evaluation, we gathered all the insights as limitations and possible future avenues for our typology.
@@
  • A preliminary study on the design implications of our typology and discussion on the utility of this typology in the visualization community.
@@ 2 Background & Related Work
Decision making is a broad topic studied across many disciplines, and visualization researchers emphasize that supporting decision making is a vital goal of data visualization [25, 38, 4, 15, 5]. To develop effective support, the visualization community utilizes typologies as tools to reason about visualization designs and explain problems and user tasks addressed by visualization. We review the relevant literature from these areas and discuss the significance to this work.
@@ 2.1 Decision Making & Visualization
The visualization community tends to focus on decision making from three broad lenses: cognition and perception, decision process support, and domain-specific decision support systems. In a cognitive context, decision-making is often framed in terms of rationality and judgement[54]. In this view, the most rational decisions are ones that maximize expectation of a function[29]. However, a number of human decisions are not rational[64], a fact that has impacted the research proclivities of the visualization community in two significant ways. The first is an increase in research on computational decision-making support [68, 38, 77]. While promising, it is unclear and often difficult to determine under what conditions users will accept help from or implement decisions of computational models [52, 31]. The second impact is the increased study of cognitive bias and perception in visualization contexts. Cognitive biases explain irrational decision-making as the result of using heuristic functions to evaluate information in a decision-making cognitive process[24].
@@ 2.2 Visualization Typologies
Two of the most prominent and influential visualization task typologies in the community are the analytical task typology by Amar et al. [3] and the multi-level task typologies by Brehmer and Munzner [47, 13]. The analytical tasks typology addressed a gap in highly data-centric typologies by taking an analytical-centric approach. This typology consists of ten tasks that describe “primitive" analyses of data, or low-level tasks [3]. The multi-level task typology further expanded upon these tasks to incorporate higher levels like “why?", “how?", and “what", which can be composed from lower-level tasks [47, 13]. However, these tasks are not suitable for describing decision-making in compact yet meaningful ways. Consider the decision of buying a car by choosing among a set of available used cars. This task, in its current form, cannot yet be represented in either the analytical or multi-level task typologies because more information is needed on the analytical activities and data. In other words, these typologies are oriented towards describing potential solutions or approaches for solving a problem, they do not describe the decision problem or how decisions are made. When we can describe decision problems in abstract yet precise ways, we can begin to reason about solutions more holistically.
@@ 3 Methodology
This section describes our methodology for developing a typology of decision-making tasks. Our approach involved gathering design goals from a literature review of publications in the following fields: visualization relating to decision-support tools, cognitive psychology relating to mental processes for decision-making, and other scientific literature relating to decision-making. Specifically, we iteratively analyzed design studies and visual analytics application papers to extract example scenarios and relevant contexts relating to decision-making, as detailed in subsections 3.1 and 3.2. This analysis led to a set of design goals we believe our typology should fulfill in order to accurately represent the decision processes described in the corpus’s case studies and decision-making scenarios. These design goals are described in section 3.3. Next, we shaped the typology using theory frameworks for decision-making by keeping in mind the goals we derived from the application papers, as explained in section 3.4. The result of this design process is our proposed typology of decision-making tasks for visualization, which we present in detail in Section 4.
3.1 Application Survey Corpus
@@
To refine our selection, we prioritized papers featuring specific domain applications, and eliminated generic systems lacking evaluations with domain experts. This guaranteed that the chosen tools were utilized by domain users with actual decisions and requirements that could be validated. For instance, papers solely based on hypothetical case studies were excluded (e.g., Podium[67], Zooids[44]). This filtering process resulted in a final corpus of 69 papers that directly contribute to our study’s objectives, and is available in the Supplemental Materials.
@@ 3.2 Coding Process
Our coding process followed an inductive approach to derive the design goals for our typology from the survey corpus outlined in the preceding section. The process involved a team-based approach with two coders, who are the first two authors of this paper. For each paper, they determined the overall decision goal by analyzing the task abstraction, requirements, and evaluation sections, and other supplementary materials, such as figures, video demonstrations, and supplementary materials (if available). The identified decision goals of the tools informed the design goals for a task typology. In addition, we noted unique decision contexts, such as sub-decisions that supported the overall decision, and the information that decisions use and produce. The result of this coding process resulted in seven design goals for our typology, which are discussed in the next subsection.
@@ 3.3 Design Goals for a Decision Task Typology
Through multiple brainstorming sessions involving all authors, we distilled the following list of goals that we believe a decision-making typology for visualization should fulfill in order to accurately represent the decision processes and scenarios found in the corpus. In the supplemental materials, we include excerpts from the corpus of papers that support the choice of each goal.
  1. G1
@@
  2. G2
@@
  4. G4
    Information Processing-Focused: The decision tasks within our typology should demonstrate flexibility and impartiality towards input and output types, instead emphasizing how information is processed and evaluated.
@@
3.4 Designing the Typology
Our design process began with an initial set of decision-making tasks inspired by relevant decision-making theory literature, particularly works by Fischhoff et al.[30, 29] and Simon’s decision-making phases [62, 14]. Then, using the goals from the previous subsection and other theory found in cognitive science literature[70, 54, 64, 24], we iteratively added, removed and modified the decision tasks and their properties until we achieved a stable set of decision-making tasks and definitions, resulting in the finalized typology described in Section 4. Below are key highlights of our process.
3.4.1 Types of Decision Tasks
Previous works often describe a single decision task, decide [24, 25] or choice [70, 23, 51], both referring to the action or process of deciding or resolving something. While useful for encapsulating decisions as a whole, they lack descriptiveness regarding the type of decision, that is, how a decision is made, limiting the analysis of decision-making in visualization[24].
There are two prominent works that are relevant to consider which outline possible decision tasks. We tested each of the decision tasks outlined in these works by applying them to our corpus. For example, we conducted an initial pass over the decision-making support tools and labeled them with one of the decision tasks, attempting to represent the ultimate decision the tool supports. For each categorization, we noted which of our design goals they did and did not meet.
First, Fischhoff et al.[30] describe three distinct decisions in the form of questions: “Is it time to act?", “Which is best?", and “What is possible?". While these better describe types of decisions in a general and domain-agnostic way (G5), the associated definitions are not directly applicable to visualization because they lack specificity on the use of information to reach conclusions or achieve decision goals (G4), which we found distinguished decisions and visualization designs in our corpus. Additionally, as defined by Fischhoff et al., these tasks are not composable (G1-3), which limits our ability to describe complex decision-making processes in sufficient detail for designing decision-support systems (G5-7).
The second prominent work outlining decision types is by Herbert Simon, which describes decision-making as a process comprising three stages: intelligence, design, and choice. The intelligence stage involves gathering information and identifying problems or opportunities that require a decision, the design stage involves developing, analyzing, and evaluating options, and the choice stage involves selecting and implementing one of the options, considering constraints and objectives. According to Simon, each of these stages can be further broken down into sub-stages of the same type. While Simon’s model meets goals G1-2, G5, it does not specify how information is processed within these stages (G4) or how the stages interrelate to achieve the final decision (G3).
Fischhoff and Simon’s decision types meet some of our typology design goals, but both fail to delineate differences in information processing for decisions, and neither discuss the composition of decisions in sufficient detail for visualization. Knowing the specifics of how information is processed, and therefore how decisions are made, and how smaller decisions compose together to reach a final decision, allows designers to create visualizations that better support those processes. We expand on past work by explicitly addressing how information is processed in our tasks (G1-4), and ensure the aforementioned properties exist in our typology, which we discuss next.